How do you solve for the mid point for the following points (-3,-3) and (9,-15). Please provide detail step by step instructions.
Found 2 solutions by stanbon, Mathtut:
Answer by stanbon(48558)

How do you solve for the mid point for the following points (-3,-3) and (9,-15).
------------------------
Each coordinate of the midpoint is the average of the corresponding
values in the end points:
------------------
x-coordinate: (-3+9)/2 = 6/2 = 3
y-coordinate: (-3-15)/2 = -9
------------
Midpoint: (3,-9)
======================
Cheers,
Stan H.

Answer by Mathtut(3670)

the midpoint would be sum of the x coordinates divided by 2 for the x coordinate
:
and sum of the y coordinates divided by 2 for the y coordinate
:
((-3+9)/2,(-3-15)/2)
:
(3,-9) is the midpoint of these two points